As part of an interview for a summer job with the
Answer #1

the two ship stuck together after collision so its the case of inelastic collision

50000 ton = 45359237 kg

70000 ton = 63502932 kg

20 knot = 10.2889 m/s

10 knot = 5.14444 m/s

by conservation of momentum

initial horizontal momentum = final horizontal momentum

45359237 * 10.2889 = (45359237 + 63502932) * v_x

v_x = 4.287 m/s

initial vertical momentum = final vertical momentum

63502932 * 5.14444 = (45359237 + 63502932) * v_y

v_y = 3 m/s

resultant velocity = sqrt(v_x^2 + v_y^2)

resultant velocity = sqrt(4.287^2 + 3^2)

velocity just after collision = 5.232 m/s